OpenStudy (anonymous):

y + 1.2y +1.2z 2.4yz 1.3y + 1.2z *1.2y2 + 1.2z 2.2y + 1.2z *= it is 1.2y with a little 2 up and to the side of it.

OpenStudy (anonymous):

Again, please show how you got the answer! :D Thanks!

OpenStudy (jennychan12):

2.2y + 1.2z you just add the common terms since you have both (1)y and 1.2 y same as y(1+1.2) = 2.2y and since there's no z term, it's just 1.2z so you get 2.2y + 1.2z

OpenStudy (mathstudent55):

You can add numbers with y together, and you can add numbers with z together, but you can't add numbers with y and z together. y + 1.2y + 1.2z means 1y + 1.2y + 1.2z You can add the 1y and 1.2y together. That is 2.2y You can't combine the 1.2z with the y's because y and z are different. So you have: y + 1.2y + 1.2z = 2.2y + 1.2z

OpenStudy (anonymous):

Okay thanks--all of you guys! @mathstudent55 that is what my lesson says--about you can add y+y, z+z, but not y+z. I think I am going to go look back in my lesson. I have a habit of "rushing things"! ;) Hey--I have another question. Is this going to be hard for a twelve year old? I'm 12.

OpenStudy (jennychan12):

and you're in either algebra or pre algebra?

OpenStudy (mathstudent55):

This algebra. This kind of algebra is not too hard for a 12-year old.

OpenStudy (anonymous):

Ok. Good! :) LOL--it is not any easier for me--because it takes me about an hour for every math lesson I do-and I have a BUNCH (like 7) of lessons to do! Haha-my dad is really really good at math--guess that trait did not pass down to me! :P

OpenStudy (anonymous):

@jennychan12 I am in pre-algebra. At least that is what my course says. I am also in 7th grade, GT.

OpenStudy (mathstudent55):

Let me give you an explanation of why y's and z's don't add together. It's because they stand for possibly different numbers. Let me explain with fruit instead of numbers. Let's say you have 2 apples and 3 oranges, so you have 2 apples + 3 oranges Now you are given another 4 oranges. What do you have now? 2 apples + 3 oranges + 4 oranges What do you have now in total? 2 apples + 7 oranges You wouldn't say you have 9 apples or 9 oranges, would you? You can't add apples and oranges. The same is true of variables. A variable stands for an unknown number. Let's say you have x stand for one unknown number, and y stand for another unknown number. x + y cannot be simplified becasue you don't know what x or y are. If I tell you you have 2x and now you're going to add to it 5x, 2x + 5x you can do that. 2 times an unknown plus 5 times the same unknown is 7 times that unknown, 2x + 5x = 7x

OpenStudy (anonymous):

Wow that actually made SENSE in my brain! Thanks so much! Now it is not so much a jumbled mess! :) I am watching my link to a video to teach me how to add like terms and etc. I do online school by the way. I think I have a basic grasp on what the heck I am doing now! :) :) Thank you so much for taking the time out to help me!! :D :D :D :D :D :D :D :D

OpenStudy (mathstudent55):

You're very welcome.

OpenStudy (anonymous):

:) Ok i am going to ask another Q since you are still here! :)

OpenStudy (anonymous):

6r + r – 5r 2r 1r + r 0r 7r – 5r Would it be answer choice B?

OpenStudy (mathstudent55):

r is the same as 1r. 6r + r - 5r = 6r + 1r - 5r You have 6 r's you add to that 1 r and you subtract 2 r's, how many r's do you have now?

OpenStudy (anonymous):

2r OK-I am writing that down in my notebook that r is the same as 1r.. thanks...again!!!! ;)

OpenStudy (mathstudent55):

right

OpenStudy (mathstudent55):

If you have one single r, there is no need to write it as 1r. Mathematicians are lazy. They want to do the least amount of work possible. 1r and r mean the same, so they decided to just write r, the least possible writing. Keep in mind that it's as if the 1 is there, so r + r = 2r because r + r is the same as 1r + 1r. 1r + 1r = 2r, just like 1 apple + 1 apple = 2 apples

OpenStudy (anonymous):

Ok thanks! :) :) :) Ok-2 more! I promise! Super sorry to keep pestering you with all these Q! 5x + 2(x + 6) 7x + 6 *7x2 + 12 7x + 12 7x (x + 6) *7x has a power of 2 next to it. (or variable, or whatever math term it is lol)

OpenStudy (anonymous):

It would be answer choice D right?

OpenStudy (anonymous):

Closing this question now................